### v1

# Building An Open Architecture for AI-RAN Convergence in 6G

Abstract
The proliferation of data-intensive Artificial Intelligence (AI) applications at the network edge demands a fundamental shift in RAN design, from merely consuming AI for network optimization, to actively enabling distributed AI workloads. This paradigm shift presents a significant opportunity for network operators to monetize AI at the edge while leveraging existing infrastructure investments. To realize this vision, this article presents a novel converged O-RAN and AI-RAN architecture that unifies orchestration and management of both telecommunications and AI workloads on shared infrastructure.
The proposed architecture extends the Open RAN principles of modularity, disaggregation, and cloud-nativeness to support heterogeneous AI deployments. We introduce two key architectural innovations: (i) the AI-RAN Orchestrator, which extends the O-RAN Service Management and Orchestration (SMO) to enable integrated resource and allocation across RAN and AI workloads; and (ii) AI-RAN sites that provide distributed edge AI platforms with real-time processing capabilities. The proposed system supports flexible deployment options, allowing AI workloads to be orchestrated with specific timing requirements (real-time or batch processing) and geographic targeting.
The proposed architecture addresses the orchestration requirements for managing heterogeneous workloads at different time scales while maintaining open, standardized interfaces and multi-vendor interoperability.
This paper has been submitted to IEEE for publication. Copyright may be transferred without notice.
I Introduction
5G-Advanced and 6G networks are reshaping the role of the RAN, from a traditional communication substrate to a programmable, intelligent, “sentient”, and compute-capable platform. As data-intensive and latency-sensitive applications such as Extended Reality (XR), Integrated Sensing and Communications (ISAC), and generative Artificial Intelligence (AI) emerge, there is an opportunity to rethink RAN architecture not only as a consumer of AI for internal optimization, but also as a distributed platform for hosting AI workloads [1]. This transformation demands architectural innovation that supports the convergence of communication and AI computing within the network itself—simultaneously delivering connectivity and supporting distributed AI inference and training demands. This dual capability is a foundational principle of the AI-RAN Alliance, a new telecom industry initiative dedicated to transforming networks through pervasive AI integration [2].
This paradigm shift in telecom infrastructure demands new approaches to system architecture and management. In this paper, we design network resource orchestration and deployment workflows to support AI solutions (i) for the RAN, i.e., tasked with improving its efficiency and operations, or (ii) on the RAN, i.e., leveraging an extremely distributed compute infrastructure for inference, sensing, or training at the edge. The network infrastructure and orchestration abstraction layer supports flexibility, scalability, and interoperability in enabling innovative AI-native wireless systems for next-generation applications. At the same time, to enable innovation, this network abstraction and the underlying infrastructure need to embrace openness, standardized interfaces, disaggregated components and support for multi-vendor deployments.
Figure 1: Architecture enabling AI and RAN coexistence, integrated within the proposed extension of the O-RAN framework.
Despite preliminary efforts in edge AI deployment, the telecom ecosystem currently lacks a cohesive architectural framework that can effectively manage the complex interplay between AI workloads and RAN operations at scale. The O-RAN ALLIANCE is defining open, interoperable interfaces and introducing key components such as the Service Management and Orchestration (SMO) and the RAN Intelligent Controllers (RICs), which enable intelligent control across multiple latency domains [3]. While O-RAN leverages intelligent control loops to optimize RAN performance, its architecture does not support generic AI workloads through a shared RAN infrastructure. As discussed above, this is currently being considered by the AI-RAN Alliance, which is defining coexistence mechanisms across different workloads. Recent literature has also addressed AI and RAN orchestration [4, 5, 1, 6, 7]. Specifically, [4, 5] discuss AI operations in the cloud, edge, and also RAN, orchestrating communications, computing, data, and AI models. [6] proposes a converged AI-and-O-RAN architecture with an end-to-end orchestrator that interacts with the Near-RT RIC its programmable logic units (xApps), thus limiting the scope of services that can be managed and deployed. [7] proposes operational workflows for AI-RAN. [1] presents a converged compute-communication platform that integrates both RAN and AI workloads on the same physical infrastructure. However, a comprehensive, end-to-end, and generalized architecture enabling AI and RAN coexistence following O-RAN management and orchestration principles is still missing.
Contributions. This paper takes a decisive step in this direction by proposing a converged O-RAN and AI-RAN architecture, where the O-RAN framework is extended to support the orchestration and management of AI workloads, as shown in Fig. 1. By building upon the foundational principles of modularity, disaggregation, and cloud-nativeness of the existing O-RAN architecture, this work demonstrates how AI integration can be achieved with minimal disruption and without increasing system complexity. The proposed design extends that the RAN compute infrastructure to the AI datacenter domain, introducing new monetization opportunities for public and private operators. Specifically, we make the following key contributions:
$\bullet$ Gap Analysis. We review AI-RAN activities and discuss the technical and operational challenges that must be addressed to achieve seamless AI-RAN and O-RAN integration.
$\bullet$ Unified Architecture Design. We present a new comprehensive architecture that enables AI and RAN workloads to share infrastructure resources while maintaining performance isolation and service quality, shown in Fig. 1. We also identify and address the orchestration requirements for managing heterogeneous AI and RAN workloads at different time scales. Specifically, we identify and propose two key architectural components, (i) the AI-RAN Orchestrator, at the top of Fig. 1, extends the O-RAN SMO framework to provide unified resource allocation, policy management, and workload coordination across AI and RAN services; and (ii) the AI-RAN Sites, which augment the O-RAN O-Cloud concept to further support AI within a common RAN infrastructure, enabling a distributed edge AI platform with real-time processing capabilities.
$\bullet$ Flexible Orchestration. We define comprehensive procedures for deploying AI workloads on shared RAN infrastructure, introducing two distinct workflows: (i) Batch mode for non-real-time computing requests and analytical workloads that are flexible with respect to deployment site and timeline; and (ii) Real-time mode for latency-critical AI workloads that require immediate processing and response on specific AI-RAN sites.
This work establishes an architectural blueprint for next generation intelligent radio access networks that go beyond traditional connectivity boundaries, enabling simultaneous delivery of telecommunications services and distributed edge AI capabilities on shared infrastructure.
II Toward AI-RAN
The transition towards AI-native mobile networks is being driven by three complementary research tracks within the AI-RAN Alliance: AI-for-RAN, AI-on-RAN, and AI-and-RAN. This section reviews the key principles behind each AI-RAN working group. We then analyzes how these efforts align with O-RAN ALLIANCE specification activities to establish the technical foundation for our converged architecture presented in Sec. IV.
II-A AI-RAN Principles
AI-for-RAN. The first pillar, AI-for-RAN, applies AI to enhance network automation, efficiency, and adaptability [5]. AI-for-RAN enhances RAN management through data-driven decision-making, improving performance, reliability, and cost efficiency via automation and self-healing [8, 2]. While this enables intelligent, self-optimizing networks, its scope remains limited to optimizing existing RAN functions without extending AI’s role beyond network operations [2, 3].
AI-on-RAN. Conversely, AI-on-RAN enables AI and generative AI applications to execute on the RAN infrastructure. These applications may or may not be related to cellular systems, but can leverage spare compute capacity and the geographically distributed nature of the RAN infrastructure. Traditional RAN deployments are over-provisioned for peak loads, therefore, by leveraging AI during low RAN activity periods, operators can maximize infrastructure efficiency without compromising real-time cellular communication [4]. This allows operators to provide AI-as-a-Service for diverse AI-driven applications while ensuring high-performance edge computing. By embedding AI workloads within the RAN, AI-on-RAN brings AI close to end-users and thus reduces the latency and bandwidth overhead associated with sending data to centralized data centers, making it ideal for latency-sensitive applications that require real-time inference and decision making. Examples of AI-on-RAN workloads can include distributed AI training (e.g., federated learning), video processing and computer vision applications, or edge deployments of generative AI solutions (e.g., a low-latency edge instance of an Large Language Model (LLM)), among others.
AI-and-RAN. To enable infrastructure sharing across AI applications and RAN workloads, AI-and-RAN focuses on the coexistence of heterogeneous workloads on the same compute platform. Indeed, AI and RAN workloads differ in terms of reliability, availability, and performance requirements. The RAN digital signal processing represents a non-elastic workload, where compute tasks need to be executed within specific timing constraints to ensure that the overall protocol stack can function as intended (e.g., for channel coding/decoding, scheduling, among others). At the same time, processing for a specific carrier configuration (bandwidth, number of MIMO streams) requires a fixed and predictable amount of resources. AI and generative AI inference or training solutions are more elastic, provided that enough resources are allocated to load the data structures associated with the AI models.
Figure 2: Challenges toward the integration of AI and RAN within the O-RAN architecture.
AI-RAN Orchestration Requirements. Considering these requirements, the orchestration of AI-for/on-RAN requires (i) a holistic accounting of the underlying compute and networking resources, so as to dynamically track infrastructure availability; (ii) northbound Application Programming Interfaces (APIs) to AI and RAN developers, to expose available capacity with a high granularity and to enable submitting workloads to the system, thus enabling infrastructure monetization; (iii) an interface that allows operators to express requirements and intents for the AI-RAN orchestrator (e.g., priority level for RAN workloads, maximum latency for AI inference); (iv) a scheduling strategy that combines requests, intents, and available resources into actionable policies or allocations; and (v) a southbound interface that connects to the infrastructure for AI and RAN configuration and deployment, and also allows for preemption in case time-sensitive RAN workloads need to be prioritized. Figure 1 captures these requirements in the end-to-end architecture.
II-B The Role of Open RAN
The AI-RAN Alliance mission does not include specifications or standards development. As such, it operates considering technical developments, specifications and standardization efforts within 3GPP and the O-RAN ALLIANCE. The latter has driven the design and development of networked systems that abide by the principles of openness, programmability, interoperability, and intelligent closed-loop control [3]. While focusing primarily on AI-for-RAN, leveraging AI to optimize RAN performance, the O-RAN ALLIANCE architecture has introduced several elements that provide valuable opportunities to extend AI integration beyond optimization. These include a software-based, disaggregated RAN, with atomic network functions that are deployed on the O-Cloud, a virtualization platform for cellular systems; the RICs, which connect to the RAN through open interfaces and enable the onboarding of custom intelligence in xApps (for the Near-RT RIC) and rApps (for the Non-RT RIC); the SMO, which provides management and orchestration across the whole network; and pipelines and workflows for AI-for-RAN.
The SMO and Non-RT RIC concern network management in the time scale of seconds or more, for RAN services and RAN performance/efficiency, respectively. The latest O-RAN architecture decouples them in a service-based framework, where the SMO combines multiple independent services communicating through the standardized SMO Services (SMOS) interface. The O2 interface connects the SMO with the O-Cloud [9] for infrastructure management and workload deployment.
We argue that the cloud-native architecture of Open RAN can streamline the development and adoption of AI-RAN innovation, through open interfaces and service composition that allow to onboard software, workloads, and workflows from multiple vendors, providers, and users. The open ecosystem fosters multi-vendor collaboration, ensuring that AI algorithms can be developed and deployed across diverse hardware and software platforms without vendor lock-in and with well-defined procedures. In addition, building on an existing architecture makes it possible to avoid unnecessary complexity, extending O-RAN components to also manage AI workloads rather than introducing additional orchestration solutions.
Building on this foundation, the following section identifies the critical challenges facing converged open AI-RAN architectures and how to address them.
III Challenges and Open Issues Toward AI-RAN and O-RAN Convergence
Figure 2 summarizes the main challenges toward effective AI-and-RAN orchestration and integration with O-RAN systems. These challenges arise from deploying AI into O-RAN ’s open and programmable architecture, which requires robust coordination between AI and O-RAN-compliant network functions.
Architecture and Workflows Challenges. AI-RAN coexistence requires an evolution of the 3GPP and O-RAN architectures. In the control plane, the end-to-end management system needs to include an entity responsible for the coexistence between AI and RAN, which maps the functionalities of the orchestrator introduced in Fig. 1 and in Sec. II. This includes understanding what are the proper control and data exposure abstractions for such a task, what is the component of the network that can host such functionalities, and if there needs to be a hierarchical approach to orchestration.
From a workflow point of view, there is a need for data, telemetry, control, and deployment pipelines, extending the O-RAN interfaces. This includes the definition of the southbound and northbound interfaces of the orchestrator, as shown in Fig. 1 at a high level. Southbound interfaces should support the continuous collection, pre-processing, and delivery of high-quality telemetry from distributed AI-RAN components, accommodating the latency requirements of different monitoring, inference, or control applications. The design of these interfaces comes with the unique opportunity to enable open and systematic processes for the creation of large-scale datasets. Finally, what Key Performance Measurements (KPMs) are better suited to support orchestrator tasks is an open question.
AI-RAN-specific procedures must be defined within the architectural framework to manage the lifecycle of AI workloads. This implies regulating access to RAN resources, orchestrating deployment across nodes, and ensuring compliance with the operator’s policies. To do so, current O-RAN interfaces (e.g., O1, O2, E2, A1) require targeted modifications to support the dual operation of RAN control loops and AI workflows. Without these foundational elements, the AI-RAN vision cannot scale beyond isolated use cases into a fully integrated, intelligent network infrastructure.
Orchestration Challenges. From an infrastructure point of view, as shown in Fig. 2, an AI-RAN system includes a heterogeneous pool that combines compute with different AI accelerators (e.g., Graphics Processing Units (GPUs)), storage, networking elements, radios, and spectrum, which the orchestrator needs to jointly take into account when allocating resources. RAN nodes exhibit diverse processing capabilities, also based on deployment location. Differences in computational power, memory, and latency requirements can lead to inefficiencies in workload distribution [1, 10]. The heterogeneity also extends to the workloads: as discussed in Sec. II, AI and RAN often have different workload patterns and requirements. AI itself can range from lightweight inference for real-time network optimization or value-added services to high-intensity model training. Moreover, the unpredictability of AI workloads, and in particular that of large-scale training tasks, can introduce latency spikes that may negatively impact RAN performance [11, 12]. This complicates both monitoring, which needs to account for real-time variations in the workload intensity, and orchestration, which needs to combine multi-scale and proactive coexistence approaches.
Resource Allocation Challenges. The resource allocation problem for the orchestrator needs to keep into account multiple factors including operators’ priorities, the multi-dimensionality of the problem (encompassing a distributed infrastructure across different geographical locations with heterogeneous capabilities), and the potential presence of a hierarchical control structure with varying granularity, from a centralized orchestrator to distributed components. The orchestrator input also comes from different stakeholders (e.g., AI developers, RAN vendors, operators), whose orthogonal requirements may need to be harmonized and reconciled.
Use Case Challenges. The AI-RAN orchestrator enables monetization strategies for next-generation wireless. These depend on identifying trade-offs, prioritization, and provisioning strategies to satisfy both AI and RAN tenants. To successfully cater to stakeholders representing AI-on-RAN use cases, the orchestrator northbound AI APIs need to align with best practices in the AI and cloud industries. The coexistence and monetization challenge also extends to energy efficiency. With RAN energy consumption being the highest operational expenditure for operators to date, it is important to adopt an energy-efficient and opportunistic approach to the acceleration of AI. AI models are usually associated with high computational demands, in contrast to recent trends for low-power RAN components. Hardware acceleration offers potential mitigation but requires careful scheduling to ensure energy-efficient performance [11].
Integrating third-party AI workloads also presents challenges to security, privacy, and access management. The orchestrator APIs and the end-to-end control and infrastructure management pipelines need to ensure accountability and tracking of security roles and privileges. One of the primary concerns is the deployment of AI models or workloads which include vulnerabilities in the software stack that may affect the stability of the RAN. Given the critical nature of RAN operations, such attacks could lead to network congestion, degraded service quality, or even large-scale outages [13].
In the next section, we introduce an architecture designed to address these challenges while adhering to the design principle of minimizing system complexity, and considering factors such as privacy and security.
IV Converged AI-RAN and O-RAN Architecture
Building on the overall network architecture shown in Fig. 1 and the requirements identified above, Figs. 3 and  4 detail the specific AI-RAN extensions to O-RAN components and their interfaces. This converged architecture has been designed to enable two concurrent workflows for the orchestration, deployment, and execution of AI workloads, i.e., batch and real-time. In batch mode, AI tasks are submitted to the centralized orchestrator, and distributed when/where compute capacity is available. This can be used, for example, for AI training, or general AI processing that requires acceleration on GPU but does not have constraints on the deployment location and timing of the deployment. In real-time mode, AI developers submit jobs for direct execution to a specific cluster (or geographical area of interest) and the request will be accommodated by the orchestrator as long as resources are available.
Figure 3: Extension of the O-RAN SMO to support AI-RAN use cases and orchestration. The components in orange, together with the SMOS, are part of the O-RAN SMO architecture. Components in blue represent extensions to accommodate AI-RAN requirements.
Figure 3 illustrates the architectural extension of the SMO toward an AI-enabled SMO (AI-SMO). Besides O-RAN SMO components connected through the SMOS bus, a set of new micro-services covers the following functionalities: resource allocation, AI workload automation, authentication and workload validation for the AI components, and exposure of interfaces to interact with the AI-RAN elements. By embedding the AI-RAN orchestrator in the SMO, we provide the orchestrator with (i) access to granular monitoring data on the infrastructure already collected through the O-RAN O1 and O2 interfaces; and (ii) a global, centralized viewpoint on the distributed RAN infrastructure that enables a holistic approach to allocation of resources to either AI or RAN workloads. The workload validation and AI user authentication address security concerns through an authorization process that vets users and AI tasks before they are admitted to the system. It validates the identity of AI users through the northbound interface to handle requests for training or inference execution. Upon successful validation, the component issues an authorization token granting access to the appropriate resources. Then, upon availability of resources determined by the orchestrator, AI task requests are forwarded to the AI workloads automation component, which triggers their deployment to the AI-RAN infrastructure sites. This components operates in parallel to equivalent services for RAN workloads, including the Network Function Orchestration (NFO) and Federated O-Cloud Orchestration and Management (FOCOM) services, with which it coordinates through the SMOS. Batch AI and RAN workloads are deployed to the selected AI-RAN edge sites using an extension of the O2 interface (AI-O2), which also communicates to the edge sites that resource allocation schedules or policies defined by the AI-RAN Orchestration service. Quality of Service (QoS) policies for AI-O2 traffic can further help by prioritizing RAN workloads over less time-sensitive AI tasks.
Figure 4: Architecture for an edge cloud site that supports AI-RAN solutions on a compute cluster managed by the AI-O-Cloud. The top part of the figure shows the AI workloads and RAN workloads.
At the other end of the AI-O2 interface, the AI-RAN edge site, shown in Fig. 4, is responsible for the granular, container-level service orchestration. Here, the O-Cloud extends into an AI-ready version that can manage containerized workloads for AI and RAN solutions. Figure 4 also shows the taxonomy of workloads that can be executed on AI-RAN sites. These include the AI-driven Central Units (CUs) and Distributed Units (DUs), with AI-for-RAN solutions (e.g., dApps [14]) providing intelligence within the protocol stack. An additional AI-for-RAN category, outside of the RAN CU/DUs, is represented by AI-accelerated Near-RT RICs, which can be hosted at the edge to satisfy near-real-time inference and control constraints. AI-on-RAN solutions can include training for AI models, generative AI services (e.g., chatbots), sensing based on AI, and other edge inference tasks, among others. As an instance, Fig. 5 illustrates the GPU utilization for two GPU-accelerated nodes in a prototype of an AI-RAN cluster, based on [15]. The system shares two applications, a GPU-based DU (based on NVIDIA ARC), with 15 connected users exchanging up to 1.5 Gbps of downlink traffic, and a chatbot using an Ollama-based LLM. The first represents an AI-for-RAN workload, and the second is an AI-on-RAN task. In this instance, the system has enough GPU capacity to support both solutions, showing the potential of RAN infrastructure sharing.
Figure 5: GPU utilization over time for a cluster with 2 GPUs (maximum GPU utilization at 200%), with AI-for-RAN (GPU-based DU) and AI-on-RAN (LLM chatbot) dynamic workloads.
The AI-RAN site is equipped with compute (including AI accelerators, e.g., GPUs), storage, and networking resources that are presented and organized through a container orchestration platform (e.g., Kubernetes or RedHat OpenShift). Both of these components are managed by the AI-O-Cloud Infrastructure Management Service (IMS) and Deployment Management Service (DMS). These extend O-RAN services that terminate the O2 interface for infrastructure configuration and workload lifecycle management, respectively. In the AI-O-Cloud, they also translate policies and schedules determined by the SMO into AI infrastructure configurations and AI workload deployment, while enabling coexistence with the RAN.
Figure 6: Procedure for the authentication and deployment of real-time AI tasks on the AI-RAN infrastructure.
In Fig. 6, we propose a procedure for the real-time deployment of AI workloads that combines (i) pre-authentication with the AI-SMO, with an indication of computing requirements at specific AI-RAN sites; (ii) the definition of a resource sharing policy in the AI-RAN Orchestrator, which is then communicated to the edge site through AI-O2; (iii) the use of a Multi-access Edge Computing (MEC) interface, adopting the ETSI standards for edge computing, by the pre-authenticated users to submit real-time AI workloads to a specific AI-RAN site; and (iv) involvement of the DMS, which allocates resources and deploys the AI services with minimal latency, as long as the amount of resources required is compatible with the policy defined by the AI-RAN Orchestrator. If not, the AI task is not admitted, and the process resumes at the AI-SMO, where the user can explore options to increase priority of the task (e.g., through different monetization tiers) or submit it as a batch workload or to a different AI-RAN site.
V Conclusions
We presented an end-to-end network architecture for AI-RAN coexistence. After reviewing the AI-RAN Alliance activities and O-RAN architecture, we discussed the challenges associated to operating AI and RAN on the same infrastructure. We then proposed an extension of the O-RAN architecture and O-RAN components that enables the deployment and management of AI workloads on the shared RAN infrastructure, including different modes for managing AI workloads with different requirements for deployment latency and location.