### v2

# Beyond Connectivity: An Open Architecture for AI-RAN Convergence in 6G

Abstract
Data-intensive Artificial Intelligence (AI) applications at the network edge demand a fundamental shift in RAN design, from merely consuming AI for network optimization, to actively enabling distributed AI workloads. This presents a significant opportunity for network operators to monetize AI while leveraging existing infrastructure. To realize this vision, this article presents a novel converged O-RAN and AI-RAN architecture for unified orchestration and management of telecommunications and AI workloads on shared infrastructure.
The proposed architecture extends the Open RAN principles of modularity, disaggregation, and cloud-nativeness to support heterogeneous AI deployments. We introduce two key architectural innovations: (i) the AI-RAN Orchestrator, which extends the O-RAN Service Management and Orchestration (SMO) to enable integrated resource and allocation across RAN and AI workloads; and (ii) AI-RAN sites that provide distributed edge AI platforms with real-time processing capabilities. The proposed architecture enables flexible orchestration, meeting requirements for managing heterogeneous workloads at different time scales while maintaining open, standardized interfaces and multi-vendor interoperability.
This paper has been submitted to IEEE for publication. Copyright may be transferred without notice.
I Introduction
5G-Advanced and 6G networks are reshaping the role of the RAN, from a traditional communication substrate to a programmable, intelligent, “sentient”, and compute-capable platform. With emerging data-intensive and latency-sensitive use cases such as Extended Reality (XR), Integrated Sensing and Communications (ISAC), and generative AI, there is an opportunity to rethink RAN architecture not only as a consumer of AI for internal optimization, but also as a distributed platform for hosting AI workloads [1]. This transformation demands architectural innovation that supports the convergence of communication and AI computing within the network itself—simultaneously delivering connectivity and supporting distributed AI inference and training demands. This dual capability is a foundational principle of the AI-RAN Alliance, a new telecom industry initiative dedicated to transforming networks through pervasive AI integration [2].
In this paper, we design network resource orchestration and deployment workflows to enable infrastructure support for AI solutions (i) for the RAN, i.e., tasked with improving efficiency and operations; or (ii) on the RAN, i.e., leveraging the distributed compute infrastructure for inference, sensing, or training at the edge. The network infrastructure and orchestration abstraction layer supports flexibility, scalability, and interoperability to enable innovative AI-native next-generation wireless systems. At the same time, to enable innovation, this network abstraction and the underlying infrastructure need to embrace openness, standardized interfaces, disaggregated components and support for multi-vendor deployments.
Figure 1: Architecture enabling AI and RAN coexistence, integrated within the proposed extension of the O-RAN framework. The figure also includes an example of use case for AI-and-RAN coexistence. An application developer submits an AI-on-RAN workload to enable XR processing at the edge. In the meantime, a RAN vendor provides an updated AI-for-RAN slicing solution. The operator defines the policy that the infrastructure should follow to enable both services. The AI-RAN orchestrator translates the policy into resource allocation, and coordinates with the automation workflows to deploy the services in the AI-RAN sites.
In this area, the prior work focused on edge AI, open architectures, and early results on AI and RAN orchestration. Preliminary efforts in edge AI deployment [3] did not lead to a cohesive architectural framework that can effectively manage the complex interplay between AI workloads and RAN operations at scale. The O-RAN ALLIANCE introduced intelligent control loops to optimize RAN performance [4]. However, its architecture, with open, interoperable interfaces, a SMO and the RAN Intelligent Controllers (RICs), does not support generic AI workloads on a shared RAN infrastructure. As discussed above, this is instead a focus area for the AI-RAN Alliance. Recent literature has also addressed AI and RAN orchestration [5, 6, 1, 7, 8]. Specifically, [5] and [6] discuss AI operations in the cloud, edge, and also RAN, but do not propose an architecture for orchestration. [7] proposes a converged AI-and-O-RAN architecture but the end-to-end orchestrator only interacts with the Near-RT RIC programmable logic units (xApps), thus limiting the scope of services that can be managed and deployed. [8] discusses AI-RAN AI operational workflows for AI-RAN, but does not consider architectural elements. [1] presents an infrastructure design that jointly supports RAN and AI workloads, without considering AI-RAN orchestration. Therefore, a comprehensive, end-to-end, and generalized architecture enabling AI and RAN coexistence following O-RAN management and orchestration principles is still missing.
Contributions. This paper takes a decisive step in this direction by proposing a converged O-RAN and AI-RAN architecture, where the O-RAN framework is extended to support the orchestration and management of AI workloads, as shown in Fig. 1. By building upon the foundational principles of modularity, disaggregation, and cloud-nativeness of the existing O-RAN architecture, this work demonstrates how AI integration can be achieved with minimal disruption and without increasing system complexity. The proposed design extends the RAN compute infrastructure to the AI datacenter domain, introducing new monetization opportunities for public and private operators. Specifically, we make the following key contributions:
$\bullet$ Gap Analysis. We review AI-RAN activities and discuss the technical and operational challenges to achieve seamless AI-RAN and O-RAN integration.
$\bullet$ Unified Architecture Design. We present a new comprehensive architecture that enables AI and RAN workloads to share infrastructure while maintaining performance isolation and service quality, shown in Fig. 1. We also identify and address the orchestration requirements for managing heterogeneous AI and RAN workloads at different time scales. Specifically, we identify and propose two key architectural components. (i) The AI-RAN Orchestrator (Fig. 1, top) extends the O-RAN SMO framework for unified resource allocation, policy management, and workload coordination across AI and RAN services. (ii) The AI-RAN Sites augment the O-RAN O-Cloud concept to further support AI within a common RAN infrastructure, enabling a distributed edge AI platform with real-time processing capabilities.
$\bullet$ Flexible Orchestration. We define comprehensive procedures for deploying AI workloads on shared RAN infrastructure, introducing two distinct workflows for non-real-time batch and latency-critical deployments.
This work establishes an architectural blueprint for next generation intelligent radio access networks that go beyond traditional connectivity boundaries, enabling simultaneous delivery of telecommunications services and distributed edge AI capabilities on shared infrastructure.
II Toward AI-RAN
This section reviews the key elements of each AI-RAN Alliance working group (AI-for-RAN, AI-on-RAN, AI-and-RAN) and the O-RAN ALLIANCE architecture.
II-A AI-RAN Principles
AI-for-RAN. The first pillar, AI-for-RAN, applies AI to enhance network automation, efficiency, and adaptability [6]. AI-for-RAN enhances RAN management through data-driven decision-making and Digital Signal Processing (DSP), improving performance, reliability, and cost efficiency [2]. This enables intelligence and self-optimization in existing RAN functions [2, 4].
AI-on-RAN. Conversely, AI-on-RAN enables AI and generative AI applications to execute on the RAN infrastructure. These applications may or may not be related to cellular systems, but leverage spare compute capacity and the geographically distributed nature of the RAN. RAN deployments are over-provisioned for peak loads, therefore, by leveraging AI during low RAN activity periods, operators can maximize infrastructure efficiency without compromising connectivity [5]. By embedding AI workloads within the RAN, AI-on-RAN brings AI close to end-users and thus reduces the latency and bandwidth overhead for transmission to centralized data centers. This makes AI-on-RAN ideal for latency-sensitive applications that require real-time inference and decision making. It also helps distributing workloads from large data centers to the edge. Examples of AI-on-RAN workloads can include distributed AI training (e.g., federated learning), video processing and computer vision applications, or edge generative AI solutions (e.g., a low-latency edge instance of a Large Language Model (LLM), chatbots, AI agents), among others.
AI-and-RAN. To enable infrastructure sharing across AI applications and RAN workloads, AI-and-RAN focuses on the coexistence of heterogeneous workloads. Indeed, AI and RAN workloads differ in terms of reliability, availability, and performance requirements. The RAN DSP represents a non-elastic workload, where tasks need to be executed within specific timing constraints to ensure intended functionality of the protocol stack (e.g., for channel coding/decoding, scheduling). At the same time, processing for a specific configuration (bandwidth, number of MIMO streams) requires a fixed and predictable amount of resources. AI and generative AI inference or training solutions are more elastic, provided that enough resources are allocated to load the data structures associated with the AI models.
AI-RAN Orchestration Requirements. Considering these requirements, the orchestration of AI-for/on-RAN requires (i) a holistic accounting of the underlying compute and networking resources, so as to dynamically track infrastructure availability; (ii) northbound Application Programming Interfaces (APIs) to AI and RAN developers, to expose available capacity and to submit workloads to the system, thus enabling infrastructure monetization; (iii) an interface for operator requirements/intents (e.g., priority level for RAN workloads, maximum latency for AI inference); (iv) a scheduling strategy that combines requests, intents, and available resources into actionable policies or allocations; and (v) a southbound interface for AI and RAN configuration and deployment. This also allows for preemption for time-sensitive RAN workloads. Figure 1 captures these requirements in the end-to-end architecture.
II-B A Primer on Open RAN
The AI-RAN Alliance does not develop specifications or standards, but builds on and influences those within 3GPP and the O-RAN ALLIANCE. The latter has driven the design and development of networked systems that abide by the principles of openness, programmability, interoperability, and intelligent closed-loop control. A complete overview can be found in [4].
The O-RAN architecture is based on a softwarized, disaggregated RAN, with atomic network functions that are deployed on the O-Cloud, a virtualization platform for cellular systems. Through open interfaces, the RICs connect the RAN to custom intelligence in xApps (for the Near-RT RIC) and rApps (for the Non-RT RIC). The SMO provides management and orchestration across the whole network. The O-RAN architecture already features pipelines and workflows for AI-for-RAN.
The SMO and Non-RT RIC concern network management in the time scale of seconds or more, for RAN services and RAN performance/efficiency, respectively. The latest O-RAN architecture decouples them in a service-based framework, where the SMO combines multiple independent services communicating through the standardized SMO Services (SMOS) interface. The O2 interface connects the SMO with the O-Cloud for infrastructure management and workload deployment [9].
III Challenges and Open Issues Toward AI-RAN and O-RAN Convergence
The cloud-native architecture of Open RAN can streamline AI-RAN innovation. Open interfaces and service composition allow onboarding software, workloads, and workflows from multiple vendors, providers, and users. The open ecosystem fosters multi-vendor collaboration, ensuring that AI algorithms can be developed and deployed across diverse hardware and software platforms without vendor lock-in and with well-defined procedures. In addition, building on an existing architecture makes it possible to avoid unnecessary complexity, extending O-RAN components to also manage AI workloads rather than introducing additional orchestration solutions.
Effective AI-and-RAN orchestration and integration with O-RAN system, however, requires addressing several challenges, which we discuss next.
Architecture and Orchestration Workflows Challenges. AI-RAN coexistence requires an evolution of the 3GPP and O-RAN architectures. The end-to-end management system requires an entity responsible for AI-and-RAN coexistence, i.e., the orchestrator introduced in Fig. 1 and in Sec. II. Challenges include understanding what are effective control and data exposure abstractions, which is network host for such functionalities, and if orchestration requires a hierarchical approach.
From a workflow point of view, there is a need for telemetry, control, and deployment pipelines, extending the O-RAN interfaces to define the southbound and northbound interfaces of the orchestrator (Fig. 1). Southbound interfaces should support continuous telemetry collection, preprocessing, and delivery from distributed AI-RAN components, accommodating latency requirements for monitoring, inference, or control applications. This unlocks the unique opportunity to systematically create large-scale datasets. Finally, what Key Performance Measurements (KPMs) are better suited to support orchestrator tasks is an open question.
Finally, the definition of AI-RAN-specific procedures for AI lifecycle management implies regulating access to RAN resources, orchestrating deployment across nodes, and ensuring compliance with the operator’s policies. To do so, current O-RAN interfaces (e.g., O1, O2, E2, A1) require targeted modifications to support the dual operation of RAN control loops and AI workflows.
Orchestration Challenges. An AI-RAN system (as shown in Fig. 1) combines compute with different AI accelerators (e.g., Graphics Processing Units (GPUs)), storage, networking elements, radios, and spectrum. The orchestrator needs to account for such heterogeneity when allocating resources. Additional compute can be available in the core network, which however does not have the same distributed nature as the RAN. RAN nodes exhibit diverse processing capabilities, also based on deployment location, which can lead to inefficiencies in workload distribution [1, 10]. The heterogeneity also extends to the workloads. As discussed in Sec. II, AI and RAN often have different workload patterns and requirements. AI itself can range from lightweight inference for real-time network optimization or value-added services to high-intensity model training. Moreover, the unpredictability of AI workloads, and in particular that of large-scale training tasks, can introduce latency spikes that may negatively impact RAN performance [11, 12]. This complicates both monitoring, which needs to account for real-time variations in the workload intensity, and orchestration, which needs to combine multi-scale and proactive coexistence approaches. Finally, orchestration workflows can extend beyond the RAN, toward core network functions that contribute to the AI operations.
Resource Allocation Challenges. The resource allocation problem must consider operators’ priorities, the multi-dimensionality of the problem (encompassing a distributed infrastructure across different geographical locations with heterogeneous capabilities), and potentially a hierarchical control structure with varying granularity, from a centralized orchestrator to distributed components. The orchestrator input also comes from different stakeholders (e.g., AI developers, RAN vendors, operators), whose orthogonal requirements may need to be harmonized and reconciled. Moreover, operators may impose different Service Level Agreement (SLA) to AI and RAN. These may change across sites and time and need to be accounted for when allocating resources. Therefore, while solutions can be inspired by prior work on scheduling and slicing, they also need to take into account the complexity of the input, output, and dimensionality of the infrastructure partitioning problem.
Use Case, Monetization, and Energy Efficiency Challenges. The AI-RAN orchestrator enables monetization strategies for next-generation wireless by identifying trade-offs, prioritizing, and provisioning strategies to satisfy both AI and RAN tenants. To successfully cater to AI-on-RAN stakeholders, the orchestrator needs to align with AI and cloud best practices.
The coexistence and monetization challenge also extends to energy efficiency. With RAN energy consumption being the highest operational expenditure for operators to date, it is important to adopt an energy-efficient and opportunistic approach to the acceleration of AI, and to account for the impact of energy in the total costs of the AI-RAN infrastructure. AI models are usually associated with high computational demands, in contrast to recent trends for low-power RAN components. Hardware acceleration offers potential mitigation but requires careful scheduling to ensure energy efficiency [11] while accounting for non-linearity and multi-dimensionality of the AI accelerators power consumption as a function of processor load, memory utilization, and data transfer.
Integrating third-party AI workloads also presents challenges to security, privacy, and access management. The orchestrator APIs and the end-to-end control and infrastructure management pipelines need to ensure accountability and tracking of security roles and privileges. Deploying AI models or workloads with vulnerabilities in the software stack could affect the stability of the RAN. Such attacks could lead to network congestion, degraded service quality, or large-scale outages [13].
In the next section, we introduce an architecture designed to address these challenges while adhering to the design principle of minimizing system complexity, and considering factors such as privacy and security.
IV Converged AI-RAN and O-RAN Architecture
Building on the architecture in Fig. 1 and the requirements identified above, Figs. 2 and 3 detail the specific AI-RAN extensions to O-RAN components and their interfaces. This converged architecture has been designed to enable two concurrent workflows for the orchestration, deployment, and execution of AI workloads, i.e., batch and real-time. In batch mode, AI tasks are submitted to the centralized orchestrator, and distributed when/where compute capacity is available. This can be used, for example, for AI training, or general AI processing that requires acceleration on GPU but does not have constraints on the deployment location and timing of the deployment. In real-time mode, AI developers submit jobs for direct execution to a specific cluster (or geographical area of interest) and the request will be accommodated by the orchestrator according to resource availability.
Figure 2: Extension of the O-RAN SMO to support AI-RAN orchestration. The components in light blue, together with the SMOS, are part of the O-RAN SMO architecture. Components in dark blue represent extensions to accommodate AI-RAN requirements.
IV-A AI-SMO and AI-RAN Site
AI-SMO. Figure 2 illustrates the architectural extension of the SMO toward an AI-enabled SMO (AI-SMO). Besides O-RAN SMO components connected through the SMOS bus, a set of new micro-services covers the following functionalities: resource allocation, AI workload automation, authentication and workload validation for the AI components, and exposure of interfaces to interact with the AI-RAN elements. By embedding the AI-RAN orchestrator in the SMO, we provide the orchestrator with (i) access to granular monitoring data on the infrastructure already collected through the O-RAN O1 and O2 interfaces; and (ii) a global, centralized viewpoint on the distributed RAN infrastructure that enables a holistic approach to allocation of resources to either AI or RAN workloads. The workload validation and AI user authentication provide access control with an authorization process that vets users and AI tasks before they are admitted to the system. It validates the identity of AI users through the northbound interface to handle requests for training or inference execution. Upon successful validation, the service issues an authorization token granting access to the appropriate resources. The authentication mechanism can be tailored to the specific deployment operational constraints, thanks to the disaggregated design of the AI-SMO. Then, upon availability of resources determined by the orchestrator, AI task requests are forwarded to the AI workloads automation component, which triggers their deployment to the AI-RAN infrastructure sites. This component operates in parallel to equivalent services for RAN workloads, including the Network Function Orchestration (NFO) and Federated O-Cloud Orchestration and Management (FOCOM) services, with which it coordinates through the SMOS. Batch AI and RAN workloads are deployed to the selected AI-RAN edge sites using an extension of the O2 interface (AI-O2), which also communicates to the edge sites that resource allocation schedules or policies defined by the AI-RAN Orchestration service. Quality of Service (QoS) policies for AI-O2 traffic can further help by prioritizing RAN workloads over less time-sensitive AI tasks. Finally, operators can dynamically express geographically and temporally customized policies that determine AI and RAN sharing strategies.
Figure 3: Architecture for an edge cloud site that supports AI-RAN solutions on a compute cluster managed by the AI-O-Cloud. The figure shows the AI workloads and RAN workloads (top), the orchestration components (middle), and the infrastructure and interfaces (bottom).
AI-RAN Site. At the other end of the AI-O2 interface, the AI-RAN edge site, shown in Fig. 3, is responsible for the granular, container-level service orchestration. Here, the O-Cloud extends into an AI-ready version that can manage containerized workloads for AI and RAN solutions, and also potentially distributed core components such as a local User Plane Function (UPF). Figure 3 also shows the taxonomy of workloads that can be executed on AI-RAN sites. Besides AI-driven Central Units (CUs) and Distributed Units (DUs), AI-for-RAN solutions include, for example, dApps [14], real-time programmable components that extend the O-RAN architecture to provide intelligence within the protocol stack. An additional AI-for-RAN category, outside of the RAN CUs/DUs, is represented by AI-accelerated Near-RT RICs, which can be hosted at the edge to satisfy near-real-time inference and control constraints. AI-on-RAN solutions can include training for AI models, generative AI services (e.g., chatbots), sensing based on AI, and other edge inference tasks, among others.
The AI-RAN site is equipped with compute (including AI accelerators, e.g., GPUs), storage, and networking resources. These are managed by a container orchestration platform (e.g., Kubernetes or Red Hat OpenShift), which also provides tenant’s workload isolation via namespaces. Both of these components are managed by the AI-O-Cloud Infrastructure Management Service (IMS) and Deployment Management Service (DMS). These extend O-RAN services that terminate the O2 interface for infrastructure configuration and workload lifecycle management, respectively. In the AI-O-Cloud, they also translate policies and schedules determined by the SMO into AI infrastructure configurations and AI workload deployment, while enabling coexistence with the RAN.
IV-B Scalability and Real-time Deployment Workflow
Figure 4: Procedure for the authentication and deployment of real-time AI tasks on the AI-RAN infrastructure.
The service-based nature of the AI-SMO favors scalability, with different components that can be scaled to handle increasing workloads submissions or large number of sites. At the same time, the AI batch submission workflow and specific instances of optimization for the resource allocation may introduce delays when applied to large-scale systems. Further, specific AI-on-RAN task may require low-latency deployments. Examples include AI inference services that are dynamically offloaded by or need to follow highly mobile users in a network, e.g., connected autonomous vehicles. Public safety sensing applications may need rapid deployment in response to an emergency. For this reason, in Fig. 4, we propose a procedure for the real-time deployment of AI workloads. It combines (i) pre-authentication with the AI-SMO, with an indication of computing requirements at specific AI-RAN sites; (ii) the definition of a resource sharing policy in the AI-RAN Orchestrator, which is then communicated to the edge site through AI-O2; (iii) the use of a Multi-access Edge Computing (MEC) interface, adopting the ETSI standards for edge computing, by the pre-authenticated users to submit real-time AI workloads to a specific AI-RAN site; and (iv) involvement of the DMS, which allocates resources and deploys the AI services with minimal latency, as long as the amount of resources required is compatible with the policy defined by the AI-RAN Orchestrator. If not, the AI task is not admitted, and the process resumes at the AI-SMO, where the user can explore options to increase priority of the task (e.g., through different monetization tiers) or submit it as a batch workload or to a different AI-RAN site.
IV-C Profiling the AI-RAN Site
We profile the feasibility of the approach with a prototype of the AI-RAN site, focusing on AI and RAN coexistence and on their deployment latency. The results are based on multiple, independent experiments executed on the Northeastern University X5G testbed [15], which implements a private 5G network with multiple commercial O-RAN Radio Units (RUs) and fronthaul infrastructure. We integrate in X5G a single-node OpenShift cluster with a Gigabyte server with an x86 CPU and an NVIDIA A100 GPU. For the RAN software, we use the NVIDIA Aerial GPU-accelerated physical layer and OpenAirInterface. For AI workloads, we use different Ollama-based LLMs (processing prompts), and a Convolutional Neural Network (CNN) or a ResNet model trained live on CIFAR-100. The AI-SMO configures the AI-RAN site GPU with Multi-Instance GPU (MIG), an NVIDIA technology for resource partitioning. For these experiments, the GPU is partitioned in 3, with 40 GB for RAN, 20 GB for LLM, and 10 GB for ResNet/CNN.
CNNResNetCNNResNet0202040406060—— Llama-3.2-3B —— —— Llama-3-8B —— MAC Throughput [Mbps]DL Th / RAN-onlyDL Th / AI-RAN Coexistence05510101515CRC Error Rate [error/s]CRC Error Rate / RAN-onlyCRC Error Rate / AI-RAN Coexistence
Figure 5: Median values for throughput and CRC error rate with RAN only (solid bars) or AI-RAN coexistence with LLM prompts and CNN/ResNet (dotted bars).
Figure 5 reports the median throughput and CRC error rate at the MAC layer, thus including retransmissions, with 95% confidence intervals computed with the bootstrap method. The solid bars report the performance without AI; dotted bars are for experiments with User Equipment (UE) traffic, the CNN/ResNet training, and the LLM being prompted. The performance is consistent across all experiments, validating AI-RAN coexistence even in a single-node site.
Figure 6 illustrates the sequential deployment latency for different LLMs models and parameter sizes. We profile the workload deployment latency from when the AI-RAN site receives the request to when the service is available, assuming a sequential deployment for the LLM, followed by the RAN and ResNet. As can be seen, the RAN and ResNet are deployed in 1-4 s, with the longer time following the deployment of the largest Llama model that fits the MIG partition. LLM instances require between 5.8-34.4 s (average). This shows that complex workloads can be managed on the same infrastructure in a matter of seconds, and provides an indication of the dynamicity with which such workloads can be (re)deployed by the AI-SMO.
0551010151520202525303035354040Qwen2.5-1.5BLlama-3.2-1BLlama-3.2-3BLlama-3-8BTime [s]LLMRANCNN
Figure 6: Workload deployment latency on the AI-RAN site.
V Conclusions
We presented an end-to-end network architecture for AI-RAN coexistence. After reviewing the AI-RAN Alliance activities and O-RAN architecture, we discussed the challenges associated to operating AI and RAN on the same infrastructure. We then proposed an extension of the O-RAN architecture and O-RAN components that enables the deployment and management of AI workloads on the shared RAN infrastructure, including different modes for managing AI workloads with different requirements for deployment latency and location. Near-term extensions of this work include further development of the architecture into an end-to-end prototype serving as a platform to validate resource allocation strategies. In the medium term, it will also facilitate studying how interfaces and design can be streamlined into O-RAN and AI-RAN Alliances activities and specifications.